OpenStudy (anonymous):

Perform the Indicated operations. If the matrix does not exist, write impossible.

OpenStudy (anonymous):

\[2\left[\begin{matrix}-1 \\ 5 \\ -6\end{matrix}\right]+4\left[\begin{matrix} -3x \\ 2 \\ x\end{matrix}\right]-3\left[\begin{matrix}x-2 \\ 3 \\ 1\end{matrix}\right]\]

OpenStudy (anonymous):

I don't understand what to do.

OpenStudy (anonymous):

nyan cat cant do matrix problems

OpenStudy (anonymous):

the number in front of a matrix means multiply everything in matrix by that number. To add two matrices they must have the same dimensions (here all 3 x 1) , and you add the same places together.

OpenStudy (anonymous):

what about the x values?

OpenStudy (anonymous):

You treat x like any other number. For Example, the bottom row should give you: \[2\cdot (-6)+4\cdot x-3\cdot 1=-15+4x\] and that is all you can do, unless you have some other information telling you what x is...